princeevee:

Which equation represents a parabola that has a focus of (0, 0) and a directrix of y = 2? x2=−4(y−1) x2=−4y x2=−y x2=−(y−1)

princeevee:

@Hero

Hero:

One moment while I switch from tablet tablet to laptop

princeevee:

...

Hero:

Sorry for taking so long. I just got off from work. You caught me in a transition stage.

Hero:

Anyway you can use the distance formula to find the equation of the parabola.

Hero:

The distance formula for finding the equation of a parabola is \[(x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\] Where the focus is \((x_1, y_1) = (0,0)\) and the directrix is \((x_2, y_2) = (x, 2)\)

Hero:

Notice that plugging the appropriate points in to the formula yields: \((x - 0)^2 + (y - 0)^2 = (x - x)^2 + (y - 2)^2\) @princeevee would you mind attempting to simplify the above?

princeevee:

hmmm...

princeevee:

(X)^2 + x^2 + y^2?

Hero:

Not quite. Let's take it one step at a time. What does \((x - 0)^2\) simplify to?

Hero:

@princeevee are you there?

princeevee:

sorry, connection machine broke

Hero:

Is it fixed now?

Hero:

When you simplify the equation after inputting the values, you should have ended up with \(x^2 + y^2 = (y - 2)^2\) The next step is to expand the \((y - 2)^2\) on the right side of the equation.

princeevee:

crap....

Hero:

@princeevee are you able to expand \((y - 2)^2\) ?

princeevee:

its kinda a bit difficult for me...can you do it?

princeevee:

@Ultrilliam

princeevee:

just for some extra help

Hero:

I'll show you how to expand \((y - 3)^2\): \begin{align*}(y - 3)^2 &= (y - 3)(y - 3) \\&=y(y - 3) - 3(y - 3) \\&=y^2 -3y - 3y + 9 \\&=y^2 - 6y + 9\end{align*}